A sample of college students was asked how much they spent daily on a cell phone plan (to the nearest peso).

Monthly Cell Phone Plan Cost (P)

Number of Students

10 - 18

6

19 - 27

8

28 - 36

10

37 - 45

7

what is the total frequency


1
Expert's answer
2021-12-17T07:55:08-0500
"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c:c}
 Classes & Lower & Upper & M_i & f_i & M_i\cdot f_i \\ \hline
 10-18 & 10 & 18 & 14 & 6 & 84 \\
 \hdashline
 19-27 & 19 & 27 & 23 & 8 & 184 \\
 \hdashline
 28-36 & 28 & 36 & 32 & 10 & 320 \\
 \hdashline
 37-45 & 37 & 45 & 41 & 7 & 287 \\
 \hdashline
 & & & Sum= & 31 & 875 \\
 \hdashline
\end{array}"

Total frequency


"\sum_if_i=6+8+10+7=31"

"mean=\bar{x}=\dfrac{1}{n}\sum_i(M_i\cdot f_i)"

"=\dfrac{1}{31}(14(6)+23(8)+31(10)+39(7))"

"=\dfrac{875}{31}\approx28.2258"

Total frequency is "31."

The sample mean for these grouped data is "\bar{x}=\dfrac{875}{31}\approx28.2258."